# Two-dimensional inference of divertor plasma characteristics: advancements to a multi-instrument Bayesian analysis system

### v1

@@ Abstract
An integrated data analysis system based on Bayesian inference has been developed for application to data from multiple diagnostics over the two-dimensional cross-section of tokamak divertors. Tests of the divertor multi-instrument Bayesian analysis system (D-MIBAS) on a synthetic data set (including realistic experimental uncertainties) generated from SOLPS-ITER predictions of the MAST-U divertor have been performed. The resulting inference was within 6%, 5% and 30% median absolute percentage error of the SOLPS-predicted electron temperature, electron density and neutral atomic hydrogen density, respectively, across a two-dimensional poloidal cross-section of the MAST-U Super-X outer divertor.
To accommodate molecular contributions to Balmer emission, an advanced emission model has been developed which is shown to be crucial for inference accuracy. Our D-MIBAS system utilises a mesh aligned to poloidal magnetic flux-surfaces, throughout the divertor, with plasma parameters assigned to each mesh vertex and collectively considered in the inference. This allowed comprehensive forward models to multiple diagnostics and the inclusion of expected physics. This is shown to be important for inference precision when including molecular contributions to Balmer emission. These developments pave the way for accurate two-dimensional electron temperature, electron density and neutral atomic hydrogen density inferences for MAST-U divertor experimental data for the first time.
@@
Keywords: Bayesian inference, integrated data analysis, divertor physics, Balmer line emission, plasma-molecular interaction
@@ 1 Introduction
Mitigating the intense heat flux from the core plasma to the plasma-facing components within tolerable engineering limits is a crucial challenge for the realisation of fusion energy [1, 2]. This heat flux is diverted, in most current and planned magnetic confinement fusion reactor designs, towards a designated power exhaust region, the divertor. Within the divertor, complex two-dimensional (and occasionally three-dimensional) phenomena occur including: plasma-neutral interactions [3, 4]; impurity radiation [5, 6, 7]; and plasma-molecular chemistry [8, 9, 10, 11]. These interactions can lead to detachment of the divertor plasma from the divertor surface (target) by simultaneously inducing losses of power, momentum and particles (i.e, ion losses) resulting in multiple-orders-of-magnitude heat flux reductions to the divertor targets [4, 12].
@@
Our understanding of divertor physics will inform the design of the power exhaust for future, reactor-class devices [1, 13, 14, 15]. Novel diagnostic and analysis techniques that measure or infer plasma characteristics over the entire 2D cross-section of the divertor plasma would, ideally, be developed to enhance current understanding of divertor physics and to provide detailed validation of plasma-edge simulations [16] that are used to extrapolate current knowledge to planned devices. This requires either directly measuring or inferring the underlying plasma parameters such as electron temperature ($T_{\mathrm{e}}$), electron density ($n_{\mathrm{e}}$) and neutral atomic hydrogen density ($n_{\mathrm{H}}$).
@@
Current procedures for extracting 2D profiles of $T_{\mathrm{e}}$ and $n_{\mathrm{H}}$ with sufficient spatial resolution ($\sim$1 cm [17]) are limited. Point-like measurements of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ provided by Langmuir Probes and Thomson scattering diagnostics have been propagated to 2D profiles via strike point sweeping (similar to [18]). However, the constant plasma conditions and accurate equilibrium reconstruction required for this analysis limits its use and introduces significant uncertainty. Alternative approaches at JET [19] and TCV [20, 21] have made use of multi-wavelength imaging camera systems of multiple hydrogen Balmer and helium lines. They utilise atomic data from the ADAS database [22] to link the measured line-brightness to $T_{\mathrm{e}}$, $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$. However, this is an ‘ill-posed’ problem where multiple combinations of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$ can produce the same measured brightness. This, along with the uncertainties associated with the required tomographic inversion of camera images, limits the precision of inferred $T_{\mathrm{e}}$ and $n_{\mathrm{H}}$ profiles in such approaches.
@@
Integrated data analysis (IDA) offers a route to the inference of 2D profiles of underlying plasma characteristics with low uncertainties and adequate spatial resolution by collectively considering a group of diagnostics [23]. Bayesian inference techniques offer a mechanism for integrated data analysis through the utilisation of ‘forward models’ of expected diagnostic responses based on underlying plasma parameters. With these forward models, the plasma parameters that most plausibly gave rise to the data recorded by all available diagnostics can be inferred in a rigorous statistical framework. Diagnostics that directly or indirectly measure the underlying plasma characteristics can be included; the quantities of an indirect measurement (such as filtered imaging diagnostics and spectroscopy [21, 9]) are connected to the underlying plasma parameters through the forward model. Through the use of synthetic data, [23] demonstrated that this approach can result in an accurate inference of the two-dimensional poloidal cross sections of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$. Multi-wavelength imaging of emission from multiple hydrogen Balmer lines [21, 24], enabled by novel diagnostics such as MANTIS on the TCV tokamak [20, 21] and the MWI on the MAST-U tokamak [25, 24], are particularly useful to such IDA techniques due to their coverage over the entire divertor cross-section.
@@
Recent work [9, 8, 11, 26, 21] has shown that Plasma-Molecular Interactions (PMI), ultimately resulting in excited atoms that emit atomic line emission, can have a significant impact on the total emissivity of hydrogen Balmer lines 111especially that of the Balmer-$\alpha$ (3$\to$2) line. Accommodating PMI within a Balmer line emission forward model requires parameters in addition to the $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$ used to model Balmer emission in [23] (which neglected PMI). These additional parameters can give rise to multiple conflicting inference solutions: the same data corresponding to Balmer line emission can be explained by an unspecified proportion of PMI further complicating the ill-posed problem. These conflicting solutions result in a degradation of the inference, potentially preventing its reliable use on experimental data.
@@
In the MAST-U Super-X divertor, plasma-molecular chemistry has been shown to play a dominant role in hydrogen emission and particle exhaust [9, 8, 27]. The impact of PMI on the hydrogen line emission appears to be particularly significant in strongly baffled divertors, as well as alternative divertor configurations (ADC)[9, 8, 28, 27]. ‘Neutral baffling’ encloses most of the divertor region with surfaces allowing higher neutral and molecular densities [29]. For the same conditions as more conventional divertors, ADCs may permit more deeply detached conditions where there is greater loss of momentum, energy and ions. Generating additional volume between the ionising (attached) region of the divertor plasma and the target permits even higher quantities of neutral atoms and molecules (and the accompanying particle, energy and momentum sinks) [8]. Ongoing experiments at MAST-U and TCV are focused on investigating such novel ADCs [30, 31, 32, 33]. MAST-U has been designed to operate with a baffled ‘Super-X’ configuration which is characterised by enhanced (compared to conventional configurations or other ADCs) total-flux expansion [34]; the cross-sectional area of the divertor flux tubes increases from the x-point region to the outer divertor target. This is achieved by placing the divertor target at a much larger major radius than the x-point [35], increasing the ratio between the magnetic field at the x-point and the target [34]. Consistent with recent experimental results from MAST-U [8, 27, 30, 31], an increase in total flux expansion is expected to: reduce target heat loads; reduce upstream electron densities and divertor impurity fractions required for detachment; and reduce the sensitivity of the detached region length to changes in fuelling/impurity seeding [34, 35].
@@
This work presents an improved framework that allows for accurate inferences (solutions) of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$ across the entire divertor cross-section in both attached and detached regimes. Building on previous work [23], data from multiple diagnostics are combined and interpreted together using Bayesian inference. This work features filtered imaging diagnostic measurements of both helium singlet line emission and hydrogen Balmer line emission, whilst accounting for the contribution of PMI in the hydrogen Balmer line emission. By conducting the inference on a mesh aligned to poloidal magnetic flux-surfaces, physics constraints (e.g. the expected monotonic reduction in static electron pressure along a flux tube to the target [36]) can be included in the inference as prior knowledge. These ‘priors’ greatly improved the robustness of the analysis and reduced its inferred uncertainty, easing the ill-posed nature of the inference. Synthetic tests, performed by applying synthetic diagnostics with realistic experimental uncertainties to SOLPS-ITER simulations of the MAST-U Super-X divertor, have been performed to investigate the performance of this new Bayesian multi-diagnostic inference technique.
@@ 2.1 Bayesian Inference
The many different parameter combinations result in a posterior distribution. The mode of this distribution is the maximum a posteriori (MAP) estimate which is taken to be the inferred $\theta$ (solution). Sampling from the distribution provides the uncertainty on the inferred $\theta$. In this work, the uncertainty is taken to be the 95% highest density interval (HDI)22295% HDI is the most narrow interval that contains 95% of the samples.
@@
To adequately explain the (2D) data in the divertor, complex diagnostics often necessitate forward models that demand well in excess of six model parameters. Procedures [37, 38] have been developed for inference in such high-dimensional problems. However, such approaches require careful analysis and construction of the posterior distribution (see appendix B.2) to ensure the posterior distribution is correctly characterised.
@@ 2.2 Advantages and disadvantages of cell-based versus mesh-based approaches
To avoid the need for complex analysis procedures and to reduce the computational time required, one can simplify the analysis by restricting the number of model parameters. For example, inference has been performed at TCV [21] by tomographic inversion of spectrally filtered camera images, at different hydrogen emission wavelengths obtained by MANTIS [20], based on the SART technique [39]. This transformed the camera data into emissivity “data-points” at discrete spatial locations, cells, throughout a 2D poloidal cross-section. Such ‘cell-based’ inference partitions the problem into multiple, independent inferences - one for each cell’s emissivity “data-point”. This enables a forward model of the emissivity of a cell rather than a full model of the camera data (which requires accounting for the plasma parameters in all cells simultaneously). Similar cell-based inferences have also been applied to hydrogen Balmer line emission in JET [26] and He I emission in TCV [40].
@@
    When an independent inference is made in each cell, information from neighbouring cells are not accounted for. This leads to additional limitations:
@@
      Cell-based inference reduces the ability to simultaneously include many diagnostics in the overall inference. For example, [21] showed, using a cell-based approach, that accurate estimates of neutral densities could be obtained, but only at the location of the Thomson scattering measurement (as this strongly constrains $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$). However, since each cell was treated independently, the information was lost outside of the location of the Thomson scattering measurements.
@@
Rather than many independent cells, a “mesh-based” inference simultaneously considers plasma-parameters at vertices spatially distributed over a mesh. By defining this mesh over a two-dimensional poloidal cross-section, through interpolation between vertices and the assumption of toroidal symmetry, plasma parameter values are available at all spatial locations throughout the divertor. This allows for comprehensive forward models to predict diagnostic data without the need for simplification of the data (e.g. through tomographic inversion). Crucially, such joint treatment of parameters retains information; by demanding a certain level of spatial smoothness of the plasma parameters, discrete information from Thomson scattering measurements influences all nearby vertices. Furthermore, combining diagnostics has a boot-strapping effect on improving inference uncertainties. Consider a camera pixel’s chord which passes through a Thomson scattering measurement location, the $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ information provided by the Thomson scattering constrains the emission arising from that region of the plasma which, in turn, limits the emission arising from all other regions of plasma along the chord. The cost of jointly considering plasma parameters is a drastic increase in the total number of model parameters and a more complex posterior distribution. However, previous work [23] demonstrated that the additional information held by not pre-inverting imaging data and jointly considering parameters distributed over a mesh across a poloidal cross-section allowed for inferences with high accuracy despite the additional complexities in the posterior distribution.

### v2

# Bibliography.bib

@@ Abstract
An integrated data analysis system based on Bayesian inference has been developed for application to data from multiple diagnostics over the two-dimensional cross-section of tokamak divertors. The divertor multi-instrument Bayesian analysis system (D-MIBAS) has been tested on a synthetic dataset (including realistic experimental uncertainties) generated from SOLPS-ITER predictions of the MAST-U divertor. The resulting inference was within 6%, 5%, and 30% median absolute percentage error of the SOLPS-ITER-predicted electron temperature, electron density, and neutral atomic hydrogen density, respectively, across a two-dimensional poloidal cross-section of the MAST-U Super-X outer divertor.
To accommodate molecular contributions to Balmer emission, an advanced emission model has been developed. This is shown to be crucial for inference accuracy. Our D-MIBAS system utilises a mesh aligned to poloidal magnetic flux-surfaces, throughout the divertor, with plasma parameters assigned to each mesh vertex and collectively considered in the inference. This allowed comprehensive forward models of multiple diagnostics and the inclusion of expected physics. This is shown to be important for inference precision when including molecular contributions to Balmer emission. These developments pave the way for accurate, two-dimensional electron temperature, electron density, and neutral atomic hydrogen density inferences for MAST-U divertor experimental data for the first time.
@@
Keywords: Bayesian inference, integrated data analysis, divertor physics, Balmer line emission, plasma-molecule interaction
@@ 1 Introduction
Mitigating the intense heat flux, from the core plasma to the plasma-facing components, to keep it within tolerable engineering limits is a crucial challenge for the realisation of fusion energy [Wenninger2014, Pitts2019]. This heat flux is diverted, in most current and planned magnetic confinement fusion reactor designs, towards a designated power exhaust region, the divertor. Within the divertor, complex two-dimensional (and occasionally three-dimensional) phenomena occur including: plasma-neutral interactions [Verhaegh2019, Lipschultz_1997]; impurity radiation [Henderson2018, Reimold2015, Bernert2023]; and plasma-molecular chemistry [Verhaegh_2023, Verhaegh_2022, Verhaegh2021, Verhaegh2021b]. These interactions can lead to detachment of the divertor plasma from the divertor surface (target) by simultaneously inducing losses of power, momentum, and particles (i.e., ion losses) resulting in multiple-orders-of-magnitude heat flux reductions to the divertor targets [Lipschultz_1997, Lipschultz2007a].
@@
Our understanding of divertor physics will inform the design of the power exhaust for future, reactor-class devices [Wenninger2014, Osawa2023, Hudoba2023, Kuang2020]. Novel diagnostic and analysis techniques that measure or infer plasma characteristics over the entire 2D cross-section of the divertor plasma would, ideally, be developed to enhance current understanding of divertor physics and to provide detailed validation of plasma-edge simulations [WIESEN2015480] that are used to extrapolate current knowledge to planned devices. Achieving this requires either direct measurement or inference of underlying plasma parameters such as electron temperature ($T_{\mathrm{e}}$), electron density ($n_{\mathrm{e}}$), and, neutral atomic hydrogen density ($n_{\mathrm{H}}$).
@@
Current procedures for extracting 2D profiles of $T_{\mathrm{e}}$ and $n_{\mathrm{H}}$ with sufficient spatial resolution ($\sim$1 cm [Eich_2013]) are limited. Point-like measurements of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ provided by Langmuir Probes and Thomson scattering diagnostics have been propagated to 2D profiles via strike point sweeping (similar to [Allen1997]). However, the constant plasma conditions and accurate equilibrium reconstruction required for this analysis limits its use and introduces significant uncertainty. Alternative approaches at JET [Karhunen2020] and TCV [Perek2019submitted, Perek_2022] have made use of multi-wavelength imaging camera systems of multiple hydrogen Balmer and helium lines. They utilise atomic data from the ADAS database [adas] to link the measured line-brightness to $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$. However, this is an ‘ill-posed’ problem where multiple combinations of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$ can produce the same measured brightness. This, along with the uncertainties associated with the required tomographic inversion of camera images, limits the precision of inferred $T_{\mathrm{e}}$ and $n_{\mathrm{H}}$ profiles in such approaches.
@@
Integrated data analysis (IDA) offers a route to the inference of 2D profiles of underlying plasma characteristics with low uncertainties and adequate spatial resolution by collectively considering a group of diagnostics [Bowman_2020]. Bayesian inference techniques offer a mechanism for integrated data analysis through the utilisation of ‘forward models’ of expected diagnostic responses based on underlying plasma parameters. With these forward models, the plasma parameters that most plausibly gave rise to the data recorded by all available diagnostics can be inferred in a rigorous statistical framework. Diagnostics that directly or indirectly measure the underlying plasma characteristics can be included; the quantities of an indirect measurement (such as filtered imaging diagnostics and spectroscopy [Perek_2022, Verhaegh_2022]) are connected to the underlying plasma parameters through the forward model. Through the use of synthetic data, [Bowman_2020] demonstrated that this approach can result in an accurate inference of the two-dimensional poloidal cross sections of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$. Multi-wavelength imaging of emission from multiple hydrogen Balmer lines [Perek_2022, Wijkamp2023], enabled by novel diagnostics such as MANTIS on the TCV tokamak [Perek2019submitted, Perek_2022] and the MWI on the MAST-U tokamak [Feng2021, Wijkamp2023], are particularly useful to such IDA techniques due to their coverage over the entire divertor cross-section.
@@
Recent work [Verhaegh_2023, Verhaegh_2022, Verhaegh2021b, Perek_2022, Karhunen2023] has shown that plasma-molecule interactions (PMI), ultimately resulting in excited atoms that emit atomic line emission, can have a significant impact on the total emissivity of hydrogen Balmer lines 111especially that of the Balmer-$\alpha$ (3$\to$2) line. Accommodating PMI within a Balmer line emission forward model requires parameters in addition to the $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$ used to model Balmer emission in [Bowman_2020] (which neglected PMI). These additional parameters can give rise to multiple conflicting inference solutions: the same data corresponding to Balmer line emission can be explained by an unspecified proportion of PMI, further complicating the ill-posed problem. These conflicting solutions result in a degradation of the inference, potentially preventing its reliable use on experimental data.
@@
In the MAST-U Super-X divertor, plasma-molecular chemistry has been shown to play a dominant role in hydrogen emission and particle exhaust [Verhaegh_2023, Verhaegh_2022, Verhaegh2023c]. The impact of PMI on the hydrogen line emission appears to be particularly significant in strongly baffled divertors, as well as alternative divertor configurations (ADC) [Verhaegh_2023, Verhaegh_2022, Verhaegh2023c]. ‘Neutral baffling’ encloses most of the divertor region with surfaces allowing higher neutral and molecular densities [Reimerdes2021]. For the same conditions as more conventional divertors, ADCs may permit more deeply detached conditions where there is greater loss of momentum, energy, and ions. Generating additional volume between the ionising (attached) region of the divertor plasma and the target permits even higher quantities of neutral atoms and molecules (and the accompanying particle, energy, and momentum sinks) [Verhaegh_2023]. Ongoing experiments at MAST-U and TCV are focused on investigating such novel ADCs [Moulton2023, Harrison2023, Verhaegh2023d, Theiler2017]. MAST-U has been designed to operate with a baffled ‘Super-X’ configuration which is characterised by enhanced (compared to conventional configurations or other ADCs) total-flux expansion [Lipschultz_2016]; the cross-sectional area of the divertor flux tubes increases from the x-point region to the outer divertor target. This is achieved by placing the divertor target at a much larger major radius than the x-point [Havlickova2015], increasing the ratio between the magnetic field at the x-point and the target [Lipschultz_2016]. Consistent with recent experimental results from MAST-U [Verhaegh_2023, Verhaegh2023c, Moulton2023, Harrison2023], an increase in total flux expansion is expected to: reduce target heat loads; reduce upstream electron densities and divertor impurity fractions required for detachment; and reduce the sensitivity of the detached region length to changes in fuelling/impurity seeding [Lipschultz_2016, Havlickova2015].
@@
This work presents an improved framework that allows for accurate inferences (solutions) of $T_{\mathrm{e}}$, $n_{\mathrm{e}}$, and $n_{\mathrm{H}}$ across the entire divertor cross-section in both attached and detached Super-X regimes. Building on previous work [Bowman_2020], data from multiple diagnostics are combined and interpreted together using Bayesian inference. This work features filtered imaging diagnostic measurements of both helium singlet line emission and hydrogen Balmer line emission, whilst accounting for the contribution of PMI in the hydrogen Balmer line emission. By conducting the inference on a mesh (see Section 2.2) aligned to poloidal magnetic flux-surfaces, physics constraints (e.g. the expected monotonic reduction in static electron pressure along a flux tube to the target [Stangeby1993]) can be included in the inference as prior knowledge. These ‘priors’ are found to greatly improve the robustness of the analysis and reduce its inferred uncertainty, easing the ill-posed nature of the inference. Synthetic tests, performed by applying synthetic diagnostics with realistic experimental uncertainties to SOLPS-ITER simulations of the MAST-U Super-X divertor, have been performed to investigate the performance of this new Bayesian multi-diagnostic inference technique.
@@ 2.1 Bayesian Inference
The many different parameter combinations result in a posterior distribution. The mode of this distribution is the maximum a posteriori (MAP) estimate which, in this work, is taken to be the inferred $\theta$ (solution). Sampling from the distribution provides the uncertainty on the inferred $\theta$. In this work, the uncertainty is taken to be the 95% highest density interval (HDI)22295% HDI is the most narrow interval that contains 95% of the samples.
@@
To adequately explain the (2D) data in the divertor, complex diagnostics often necessitate forward models that demand well in excess of six model parameters. Procedures [neal2011mcmc, hoffman2014no] have been developed for inference in such high-dimensional problems. However, such approaches require careful analysis and construction of the posterior distribution (see appendix B.2) to ensure the posterior distribution is correctly characterised.
@@ 2.2 Advantages and disadvantages of cell-based versus mesh-based approaches
To avoid the need for complex analysis procedures and to reduce the computational time required, one can simplify the analysis by restricting the number of model parameters. For example, at TCV, spectrally filtered camera images (at different hydrogen emission wavelengths obtained by MANTIS [Perek2019submitted]) have been tomographically inverted (based on the SART technique [Anderson1984]) to perform inference [Perek_2022]. This transformed the camera data into emissivity “data-points” at discrete spatial locations, cells, throughout a 2D poloidal cross-section. Such ‘cell-based’ inference partitions the problem into multiple, independent inferences - one for each cell’s emissivity “data-point”. This enables a forward model of the emissivity of a cell rather than a full model of the camera data (which requires accounting for the plasma parameters in all cells simultaneously). Similar cell-based inferences have also been applied to hydrogen Balmer line emission in JET [Karhunen2023] and He I emission in TCV [Linehan2023].
@@
    When an independent inference is made in each cell, information from neighbouring cells is not accounted for. This leads to additional limitations:
@@
      Cell-based inference reduces the ability to simultaneously include many diagnostics in the overall inference. For example, [Perek_2022] showed, using a cell-based approach, that accurate estimates of neutral densities could be obtained, but only at the location of the Thomson scattering measurement (as this strongly constrains $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$). However, since each cell was treated independently, the information was lost outside of the location of the Thomson scattering measurements.
@@
Rather than splitting the plasma state into many independent cells, a ‘mesh-based’ inference simultaneously considers plasma parameters at vertices spatially distributed over a mesh. By defining this mesh over a two-dimensional poloidal cross-section, through interpolation between vertices and the assumption of toroidal symmetry, plasma parameter values are available at all spatial locations throughout the divertor. This allows for comprehensive forward models to predict diagnostic data without the need for simplification of the data (e.g. through tomographic inversion). Crucially, such joint treatment of parameters retains information; by demanding a certain level of spatial smoothness of the plasma parameters, discrete information from Thomson scattering measurements influences all nearby vertices. Furthermore, combining diagnostics has a boot-strapping333i.e. information from one diagnostic can improve the information content of another diagnostic and vice versa. effect on improving inference uncertainties. Consider a camera pixel’s chord which passes through a Thomson scattering measurement location, the $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ information provided by the Thomson scattering constrains the emission arising from that region of the plasma which, in turn, limits the emission arising from all other regions of plasma along the chord. The cost of jointly considering plasma parameters is a drastic increase in the total number of model parameters and a more complex posterior distribution. However, previous work [Bowman_2020] demonstrated that the additional information held by not pre-inverting imaging data and jointly considering parameters distributed over a mesh across a poloidal cross-section allowed for inferences with high accuracy despite the additional complexities in the posterior distribution.